OpenStudy (anonymous):

Can someone help me? I WILL MEDAL & FAN The function below shows the number of students of a school who enrolled for cooking classes. Let f(x) represent the total number of students who enrolled for the classes after x years: f(x) = 11(1.35)^x The average rate of change in the number of students who enrolled for cooking classes from the first to the fifth year is ________students per year. Round your answer to the nearest whole number.

OpenStudy (anonymous):

@Directrix

OpenStudy (mathmate):

Average rate of change of a function f(x) over the interval [a,b] is given by: average rate of change =\(\large \frac{f(b)-f(a)}{b-a}\) For the given example, f(x)=\(11(1.35^x)\) which actually interprets as at year 0, there were 11 students (\(1.35^0=1\)), and student population increases at the rate of 35% annual increase over the previous year. so average rate of change = f(5)-f(1) students / year, and round to the nearest whole number of students. See also the following link for further examples and explanations. http://cims.nyu.edu/~kiryl/Precalculus/Section_2.4-Average%20Rate%20of%20Change/Average%20Rate%20of%20Change.pdf

OpenStudy (anonymous):

Thank-you !

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

for your reply what do I have to do to get my answer?

OpenStudy (mathmate):

put a=1, b=5 and evaluate the average rate of change.

OpenStudy (mathmate):

If in doubt, also look at the link I gave you.